The sum of two numbers is 20 and their difference is 4.find the numbers.

Answer by ikleyn(52786)

.
x + y = 20,
x - y =  4.

What to do next? - But of course, add these two equations (both sides). You will get

2x = 24,  hence,  x = 12.

Now y = 20 - 12 = 8 from the first equation.

Answer.  The numbers are 12 and 8.